OpenStudy (anonymous):

what is the most important thing i need to remember about uniform circular motion? i have a test tomorrow and i dont want to die. help?!

OpenStudy (anonymous):

OK ?

OpenStudy (anonymous):

tommorow came... rembember that \[a_R=\frac{v^2}{R}=\omega^2R=f^24\pi^2R=\frac{4\pi^2}{T^2}R\] and that the acceleration is always to the center